OpenStudy (anonymous):

What property is shown below? 3 × (4 × 8) = (3 × 4) × 8 Transitive Commutative Distributive Associative

OpenStudy (anonymous):

Either 1 or 3.. I think 1 tho

OpenStudy (anonymous):

Which of the following equations is true? 4 - 3 - 2 = 2 - 3 - 4 4 ÷ 3 ÷ 2 = 2 ÷ 3 ÷ 4 4 + 3 + 2 = 2 + 3 + 4 4 + 3 × 2 = 2 + 3 × 4

OpenStudy (anonymous):

IDK what that questionmarks mean.. hha

OpenStudy (anonymous):

@alanaleigh13

OpenStudy (cj49):

C